### v1

# Leveraging Content and Acoustic Representations for Efficient Speech Emotion Recognition

Abstract
Speech emotion recognition (SER), the task of identifying the expression of emotion from spoken content, is challenging due to the difficulty in extracting representations that capture emotional attributes from speech. The scarcity of large labeled datasets further complicates the challenge where large models are prone to over-fitting. In this paper, we propose CARE (Content and Acoustic Representations of Emotions), where we design a dual encoding scheme which emphasizes semantic and acoustic factors of speech. While the semantic encoder is trained with the distillation of utterance-level text representation model, the acoustic encoder is trained to predict low-level frame-wise features of the speech signal. The proposed dual encoding scheme is a base-sized model trained only on unsupervised raw speech. With a simple light-weight classification model trained on the downstream task, we show that the CARE embeddings provide effective emotion recognition on a variety of tasks. We compare the proposal with several other self-supervised models as well as recent large-language model based approaches. In these evaluations, the proposed CARE model is shown to be the best performing model based on average performance across $8$ diverse datasets. We also conduct several ablation studies to analyze the importance of various design choices.
Index Terms:
Speech-text alignment, representation learning, self-supervised learning, emotion recognition
I Introduction
Speech Emotion Recognition (SER) involves identifying the emotional state of the speaker based on acoustic characteristics. The ability to recognize emotions from speech has profound applications across various domains, such as human-computer interaction [1], social media analytics [2], customer call centres [3] and mental health monitoring systems [4]. Despite significant advancements, SER remains a challenging problem due to the complex nature of human emotions and the difficulties in succinctly capturing the human emotions from limited supervised datasets.
Traditionally, SER systems used different acoustic properties of speech signals to analyze emotion content. The importance of the pitch contour towards this task was first highlighted by Lieberman et al. [5]. Other acoustic characteristics such as energy, intensity and speaking rate were subsequently established as markers of emotion [6]. These features, proposed as part of the Interspeech para-linguistic challenges, were high dimensional  [7, 8]. The dimensionality issue was partially addressed by Eyben et al. [9], where a minimalist set of features were proposed for SER. However, the hand crafted features suffer from lack of robustness to novel emotional datasets and tasks.
In the last decade, the research efforts in SER progressed from using hand crafted features to deep learning techniques. The network architectures include convolutional neural networks (CNN) [10, 11], long short term memory (LSTM) [12] networks, and transformer models [13]. More recently, self-supervised learning (SSL) has emerged as a viable alternative to the classical approach of training individual models for different tasks. These SSL models, trained on large amounts of unlabeled data, have been shown to generalize well to a variety of downstream tasks. The popular examples include wav2vec2.0 [14], HuBERT [15] and WavLM [16], which have shown promising results on emotion recognition [17, 18].
With the advent of large language models (LLMs), recent efforts in speech representation learning attempt to encode speech using speech-to-text encoders like Whisper [19] followed by large language modeling with adapters. Examples of such efforts include SALMONN [20] and wavLLM [21]. Further, the alignment of speech and text at embedding level was attempted by SONAR [22]. However, these efforts require a large model with billions of parameters to perform tasks like emotion recognition.
In this paper, we propose the design of a novel parameter-efficient modeling framework of representing speech in a dual encoding fashion. The proposed model is termed as Content and Acoustic Representations of Emotions, where one of the encoder arms, termed semantic encoder, aligns the speech representations with the recognition transcripts encoded using the RoBERTa model [23] at sentence level, while the non-semantic arm aligns the speech representations with low-level descriptors of speech, given by the PASE+ model [24]. Both the encoder arms are learned through distillation losses, with the weights initialized using the WavLM [16] SSL model. As the emotion recognition and understanding tasks entail both the semantic and non-semantic representations, we concatenate the two encoder outputs and learn a light-weight classification head to perform emotion recognition for the downstream tasks. We show the effectiveness of the CARE approach by evaluating on $8$ speech emotion datasets as well as on a depression detection dataset. The key contributions of our work are as follows:-
  • Designing a parameter efficient SER system, termed as CARE, consisting of a semantic encoder and an acoustic encoder.
  • Performing distillation based learning of the CARE model by leveraging pre-trained SSL models of speech - WavLM [16] and PASE+ [24]) for acoustic arm, Whisper [19] and text (RoBERTa [23]) models for semantic arm.
  • Proposing a convolution layer adaptation of text-based models for representation learning of speech.
  • Experimentation on $8$ benchmark speech datasets with diverse SER tasks - achieving state-of-art (SoTA) performance on some these tasks while outperforming models of much larger size.
II Related work
II-A Audio feature extraction for SER
Speech emotion is often conveyed through various factors encoded in speech like content, pitch, prosody, stress, loudness etc [25]. Earlier works attempted to extract wide variety of acoustic descriptors for emotion recognition [9]. Deep learning based representations have been increasingly used as low-level acoustic features. Some examples in this line of work include the SincNet architecture by Ravanelli et al. [26], problem agnostic speech representations (PASE) by Pascual et al. [27] and interpretable Gaussian filters by Agrawal et al. [28]. One such architecture, called the learnable front-end for audio classification (LEAF) [29] was used by Dutta et al. [11] for speech emotion classification. In contrast, this work designs a setting where the semantic features are also represented along with low-level acoustic features for emotion recognition.
II-B Self supervision in speech
Following the application of self-supervised models for text, a large number of SSL models emerged in speech. One of the earliest such models was proposed by Schneider at al. [30]. The technique of masking out parts of the data and subsequently predicting them, also called masked language modeling (MLM), was introduced in the domain of natural language processing [31]. A similar setting for speech was adopted by Baevski et al. [14]. An iterative training procedure, called HuBERT, proposed by Hsu et al. [15], used MLM loss and led to further improvements in various downstream tasks. The approach termed WavLM [16] explored augmenting the pre-training recipe for HuBERT with a speech separation task. Recent approaches in this direction include teacher-student learning setting [32] or disentangled learning [33, 34].
II-B1 Self supervision for SER
Siriwardhana et al. [35] proposed a multimodal model for emotion recognition wherein they processed the speech signal by a vq-wav2vec model [36] followed by a pre-trained speech-BERT [37] model. Similarly, the pre-trained wav2vec model has been used in other works [38, 39]. Pepino et al. [40] showed the effectiveness of fusing the wav2vec2.0 features along with the eGeMAPS features for emotion recognition.
We describe two recent methods in this direction - Vesper by Chen et al. [41] and emotion2vec [42]. The primary objective of Vesper was to compress a pre-trained model for emotion recognition in a teacher-student mode of training, with the teacher being initialized with the WavLM-large model. Emotion2vec [42] is a method where the data2vec [32] model is considered as the teacher which is used to train a student model. In contrast to the proposed work, the prior SSL efforts for SER largely ignore the high-level semantic content information, which is also observed to contribute significantly to emotion recognition tasks [18].
II-C Speech-text aligned representations
The alignment of speech and text modalities is an increasingly active area of interest for speech representation learning. The joint pre-training of speech and text using supervised data was proposed by Bapna et al. [43]. The MAESTRO model, proposed by Chen et al. [44], explored the learning of a shared speech-text encoder and a duration model which upsampled the text embeddings to match the speech duration. In this line of work, Speech-LM, proposed by Zhang et al. [45], investigated the modeling of unpaired speech and text data. In contrast to these efforts, the CARE model attempts to jointly model textual and non-textual speech representations.
The SONAR model [22] designs a speech encoder which aligns with the encoded textual representations at the utterance level. With the rising popularity of large language models (LLMs), several recent works have developed speech encoders that are jointly learned with LLM models. Many such works use the Whisper ASR encoder [19] with LLaMA [46] LLMs. The SALMONN model by Tang et al. [20] developed speech-music encoder for broader audio understanding, while WavLLM, proposed by Hu et al. [21], used Whisper and WavLM encoders along with low-rank adaptation (LoRA) [47] of LLMs. While these LLM based efforts use speech-text aligned representations with other acoustic encoders and enable prompt based applications, the models are inherently large-sized ($8$B) with a significant burden on training and inference. In contrast, our proposed CARE system, even with its compact size of $160$M parameters, outperforms larger models in several downstream evaluations.
Figure 1: Block diagram of the proposed CARE model. An ASR system is used to generate the transcripts for the pre-training data which is passed through a pre-trained RoBERTa model to generate targets for the semantic arm. The acoustic arm of the model is trained with PASE+ features as targets. Blocks in blue indicate either frozen components or those with no learnable parameters.
III Proposed Approach
III-A Background
III-A1 WavLM
The MLM task in the speech domain, with architectures like wav2vec2.0 [14] and HuBERT [15], advocates a style of pre-training, wherein segments of the speech signal are masked out and the transformer based models are trained to predict the masked speech regions. The WavLM [16] uses the same pre-training architecture as HuBERT as well as the clustering based discrete tokenization of audio with the addition of a de-noising objective. The speech signals are randomly mixed with noise or with other speech signals to create noisy or overlapped speech. A stack of convolutional layers transform the speech data followed by $12$ ($24$ for WavLM-large) transformer layers to predict the discrete cluster labels corresponding to clean speech. This de-noising objective, along with the masked language modeling loss, results in improved performance on non-ASR tasks compared to HuBERT or wav2vec2.0 [16].
III-A2 RoBERTa
One of the significant contributions in creating a large scale language model was proposed by Devlin et al. [31]. This architecture, called bidirectional encoder representations from transformer (BERT), was trained with two objectives on a corpus of textual data, namely, predicting words masked out in a sentence and to predict whether two sentences semantically follow each other (referred to as next sentence prediction). Liu et. al [23] trained this architecture on a larger corpus of textual data without the next sentence prediction task. This pre-trained model is known as robust optimized BERT approach (RoBERTa). We use the pre-trained base version of this model with $12$ transformer layers as the text encoder.
III-B CARE Model
The block diagram of the CARE model is shown in Fig.1. Text supervision: The annotation of text transcripts for the emotional pre-training speech datasets is highly demanding. A pre-trained ASR system offers an alternate way to generate “noisy” transcripts of speech. We use an open source ASR model, Whisper-large-v3 [19], for converting speech to text transcripts. Note that, this ASR model was not trained on emotional speech, hence the word-error-rate (WER) on emotional speech datasets is higher than those on neutral speech datasets [18, 48]. The noisy text transcripts are encoded using the RoBERTa model [23] and the mean-pooled utterance features serve as the teacher signal for the semantic arm of the CARE model.
Acoustic Supervision: The non-semantic components of speech, typically captured in low-level acoustic descriptors of speech like filter-bank energies, prosody, pitch etc are also potentially rich speech representations containing information about the underlying emotion. The PASE+ model [24] contains a series of convolutional and recurrent neural network layers which encode the speech, followed by a set of workers that regress the low-level descriptors of speech, derived from the clean input signal. We use the frame-level PASE+ features of $256$ dimensions as targets for the acoustic arm of the CARE model. These features are down-sampled by a factor of $2$ to generate targets at $50$Hz.
III-B1 Model Architecture
The speech signal is initially processed by a set of convolutional layers. The stack of $6$ transformer layers acts as the common encoder arm. Both the acoustic and semantic arms are $6$-layer transformer models. They differ in the distillation targets that are used to train the respective arms. The acoustic encoder uses the PASE+ representations as targets, while the semantic encoder uses the RoBERTa representations of the ASR transcripts as targets. Further, the acoustic arm receives a frame-level loss while the semantic encoder receives an utterance level loss. The utterance level objective function for text data avoids the need for time-aligning speech and text, one of the key difficulties in training speech-text joint models [44].
For the semantic arm, we propose a novel adaptation strategy. The last $6$ transformer layers of the pre-trained RoBERTa model are used as frozen weights for the semantic arm of the CARE model. In the semantic arm, two convolutional blocks are trained, one before and one after each transformer layer. These convolutional blocks serve the purpose of adapting the speech representations to the internal representations of the RoBERTa model. The convolutional block before the transformer layer down-samples the sequence length by a factor of $3$, while the block following the transformer layer up-samples the representations by the same factor. This ensures that the frozen RoBERTa transformer layer processes sequences with lower time-resolution than the original speech. The representations from the semantic arm are average pooled to generate an utterance level representation of the speech signal.
III-B2 Loss
An utterance level loss, $L_{utt}$ and a frame-level acoustic loss, $L_{acoust}$, are employed for training the semantic and acoustic arms, respectively. Denoting the mean-pooled output from the pre-trained RoBERTa model by $\boldsymbol{y}_{text}$ and the output from the semantic arm as $\hat{\boldsymbol{y}}_{sem}$, the semantic loss is the mean square error (MSE) loss:
$L_{sem.}=\frac{1}{N}\sum_{i=1}^{N}||\boldsymbol{y}_{text}^{i}-\hat{\boldsymbol{y}}_{sem}^{i}||_{2}^{2}$ (1)
where $N$ denotes the batch size.
For the frame level loss, let $\hat{\boldsymbol{y}}_{acoust}\in R^{N\times T\times D}$ denote the output of the acoustic arm, where $N$, $T$ and $D$ denote the batch size, number of frames per utterance and the dimension of the representation, respectively. The loss is defined as:
$L_{acoust}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}||\boldsymbol{y}_{pase}^{ij}-\hat{\boldsymbol{y}}_{acoust}^{ij}||_{2}^{2}$ (2)
where $\boldsymbol{y}_{pase}$ denotes the acoustic target from the PASE model.
III-B3 Inference
For the different downstream tasks, the model is evaluated according to the paradigm proposed in SUPERB benchmark [49]. The representations from the semantic and acoustic arms are concatenated layer-wise. A convex combination of the layer representations is fed to a classification head. The only learnable parameters on the downstream task are the convex combination weights as well as the light-weight classification head, while the weights of the common, acoustic and semantic arms are kept frozen for the downstream supervised training task. This setting is also followed for other baseline models compared in this work.
IV Experiments and Results
IV-A Pre-training
The MSP-PODCAST corpus [50] is used for the task of pre-training. A total of $149,307$ speech-turns amounting to $230$ hours of emotional speech data is used. Out of the total number of samples, $80\%$ of the data is randomly chosen as the training set while the remaining $20\%$ serves as the validation set. As the pre-training in the proposed CARE model is unsupervised, only the raw audio data is required. The Whisper-large-v3 model [^1] https://huggingface.co/openai/whisper-large-v3 is used for generating the transcripts, while the pre-trained RoBERTa model [^2] https://huggingface.co/FacebookAI/roberta-base is used for encoding the text data. The common arm is initialized with first $6$ layers of the WavLM base model [^3] https://huggingface.co/microsoft/wavlm-base, while the acoustic arm is initialized with the last $6$ layers of the WavLM base model.
TABLE I: Summary of the evaluation datasets. Class balance denotes if the datasets are balanced across classes. The last column indicates if the training/test data have common speakers. For all the datasets, we perform $5$-fold evaluation.
Datasets
# Train
Utt.
# Val.
Utt.
# Test
Utt.
# Classes
Class
Bal.
Spkr.
Ind.
IEMOCAP-4 [51] $4425$ $1102$ $1102$ 4
IEMOCAP-6 [51] $5947$ $1487$ $1387$ $6$
MELD [52] $9988$ $1108$ $2610$ $7$
CMU-MOSI [53] $1188$ $325$ $686$ $2$
DAIC-WOZ [54] $6003$ $2097$ $2097$ $2$
RAVDESS-Song [55] $704$ $132$ $176$ $6$
CaFE [56] $624$ $156$ $156$ $7$
EmoDB [57] $324$ $105$ $106$ $7$
TABLE II: Results on the different downstream datasets in terms of weighted F1-score. All numbers are averaged over $5$ random initializations of the downstream network. We also show the results of the different components of CARE in this table.
Datasets PASE+ [24]
Whisper [19]+
RoBERTa [23]
Semantic Arm Acoustic Arm CARE
Params:$8$M Params:$1.6$B Params:$110$M Params:$94$M Params:$160$M
IEMOCAP-4 [51] $56.68$ $61.97$ $66.44$ $65.91$ $\mathbf{69.39}$
IEMOCAP-6 [51] $41.38$ $49.28$ $53.05$ $52.09$ $\mathbf{55.02}$
MELD [52] $35.86$ $\mathbf{49.29}$ $47.37$ $46.98$ $48.05$
CMU-MOSI [53] $50.69$ $\mathbf{75.14}$ $64.23$ $64.17$ $66.74$
DAIC-WOZ [54, 58] $66.84$ $64.04$ $66.32$ $66.89$ $\mathbf{68.49}$
RAVDESS-Song [55] $46.05$ $9.58$ $55.23$ $56.17$ $\mathbf{60.11}$
CaFE [56] $52.86$ $13.59$ $69.23$ $71.62$ $\mathbf{76.98}$
EmoDB [57] $62.59$ $14.98$ $75.42$ $78.63$ $\mathbf{83.41}$
Avg. $51.62$ $42.23$ $62.16$ $62.81$ $66.02$
IV-B Downstream tasks
A summary of the different datasets used for evaluation is mentioned in Table I.
IV-B1 IEMOCAP
The IEMOCAP dataset consists of $151$ video recordings split into 5 sessions. Each of these sessions is a conversation between a pair of subjects, one male and one female. Each recording is split into multiple utterances. There are a total of $10,039$ utterances, each of which is labeled by human annotators as belonging to one of the $10$ emotions - “angry”, “happy”, “sad”, “neutral”, “frustrated”, “excited”, “fearful”, “surprised”, “disgusted” or “other”. Keeping in line with previous works, we do a four-way classification task where we consider “angry”, “happy”, “sad”, “neutral” and “excited” categories (with “excited” and “happy” categories merged). We also have a separate setting of $6$ emotional classes [59]. The first $6$ emotion classes are considered for this setting.
IV-B2 MELD
The MELD dataset [52] is a multi-party dataset created from video clippings of the popular TV show, “Friends”. A seven way classification task is performed on this dataset, with each utterance being labeled as one of the $7$ emotions - “angry”, “sad”, “joy”, “neutral”, “fear”, “surprise” or “disgust”.
IV-B3 CMU-MOSI
The CMU-MOSI dataset [53] has a total of $2199$ utterances. Each utterance is labeled in the range of $[-3,3]$. Following previous works, we treat this as a binary classification problem with utterances having sentiment values in the range $[-3,0)$ being classified as negative sentiment and those with values in the range $[0,3]$ considered as positive sentiment. The dataset partitioning follows the prior work by Poria et. al [60].
IV-B4 DAIC-WOZ
The DAIC-WOZ dataset [54] is a benchmark dataset for depression detection, consisting of $189$ clinical interviews between the patient and the interviewer. Out of these $189$ interviews, $107$ are part of the training set while $35$ interviews are part of the development subset. The dataset suffers from a data imbalance problem, with only $30$ interviews labeled as “depressed” in the train set . In order to increase the balance, we follow [61] and extract $100$ utterances randomly from each interview, which is labeled as depressed, while only $39$ utterances are selected for interviews classified as “normal”. The utterances from each interview are chosen randomly for the $5$ splits. Following prior work [62, 63, 64, 65, 66, 61], we report results on the development set of this dataset.
IV-B5 RAVDESS-Song
The RAVDESS-Song dataset [55] has a total of $1012$ song recordings sung by $23$ different singers. Each recording in this dataset is sung in one of six different emotions, namely, “neutral”, “calm”, “happy”, “sad”, “angry” and “fear”. We conduct a speaker independent evaluation for this dataset, and create $5$ different splits. For each split, we keep recordings from $16$ singers for training, while recordings from $3$ separate singers are used for validation. The recordings from the remaining $4$ speakers are used for evaluation of the models.
IV-B6 CaFE
The CaFE dataset [56] is a Canadian French emotional dataset consisting of $936$ utterances spoken by $12$ speakers. Each utterance in this dataset is categorized as one of the seven emotions - “neutral”, “angry”, “disgust”, “sad”, “surprise”, “fear” and “happy”. Similar to RAVDESS-Song, we create $5$ speaker independent splits for this dataset. The utterances belonging to $8$ speakers are used for the training, while the remaining $4$ speakers are used for validation and testing equally. The speakers used for train, validation and test are chosen randomly for the $5$ splits.
IV-B7 EmoDB
The EmoDB dataset [57] has a total of $535$ utterances spoken by 10 different speakers. Each utterance in this dataset is categorized as one of the seven emotions - “neutral”, “angry”, “disgust”, “sad”, “boredom”, “fear” and “joy”. Similar to RAVDESS-Song and CaFE, we create $5$ different speaker independent splits for this dataset. The utterances belonging to $6$ different speakers are chosen for training while the remaining $4$ speakers are used for validation and testing equally. The speakers used for train, validation and testing are chosen randomly for each of the $5$ splits.
IV-C Loss and evaluation metrics
Since these datasets are categorical in natural, we use the cross entropy loss for training the downstream model weights. For testing, we use the weighted F1-score as the evaluation metric. Denoting the F1 score of a class $c$ with $N_{c}$ samples, by $F1_{c}$, the weighted F1-score is defined as:
$WF1=\frac{1}{\sum_{c=1}^{C}N_{c}}\sum_{c=1}^{C}N_{c}\times F1_{c}$ (3)
IV-D Implementation details
Pre-training: During pre-training, all the speech utterances from MSP-PODCAST are padded or randomly cropped to a duration of $5$ seconds. The model is trained with a learning rate of $1e$-$5$ and a batch size of $128$ with AdamW [67] as the optimizer. The model is trained for a total of $200,000$ steps and the best model parameters based on validation set performance are chosen for evaluation of downstream datasets.
Fine-tuning and evaluation: For the downstream task training, the speech signals are cropped to a maximum duration of $30$ seconds or padded to a minimum duration of $1$ second. For the depression detection dataset, DAIC-WOZ, each speech segment has a duration of $10$ seconds [61]. The representations from the layers are accumulated using a convex combination with learnable weights. The classification head is a two-layer feed-forward neural network with the ReLU activation [68]. Only the weights corresponding to convex combination of different layer representations and those in the two-layer fully connected classification head are trained for each downstream dataset, in line with the framework of SUPERB [49]. We use a batch size of $32$ with a learning rate of $1e$-$4$ and train the model for $50$ epochs. The AdamW optimizer is used here as well. All the models, including the CARE and the baseline systems, utilize the same data and classification backend, without any additional hyper-parameter tuning or model selection. Thus, the design allows the fair comparison of the representations.
IV-E Results
The results of our model across various downstream datasets are presented in Table II.
PASE+ evaluation: We extracted PASE+ features for each downstream dataset, followed by training a two-layer classification network to predict the emotion class of each utterance. The results demonstrate that acoustic arm of the proposed CARE model significantly outperforms the PASE+ baseline across all datasets.
Whisper+RoBERTa Evaluation: In this approach, transcripts are extracted using the Whisper-large-v3 model [19], and subsequently processed by a pre-trained RoBERTa model. The internal representations of the RoBERTa model are linearly combined and a two-layer classification head is trained. It is noted that, with the exception of MELD and CMU-MOSI, the semantic arm of CARE outperforms this baseline for all the other datasets. Further, the Whisper-RoBERTa system contains a large number of parameters ($10$x) compared to the proposed CARE model.
Out-of-Domain Tasks: The performance of CARE is also evaluated on three out-of-domain tasks: RAVDESS song, CaFe [56] and EmoDB [57] speech. The proposed CARE approach significantly improves over all the other models compared. The Whisper-RoBERTa model fails to generalize to these tasks, as the datasets contain song or multi-lingual speech. The results indicate that, in these instances, the acoustic arm of CARE outperforms its semantic counterpart. Despite the fact that neither component of CARE is pre-trained on datasets comprising songs or non-English speech, the model demonstrates strong generalization to these tasks. This success is attributed to the distillation based learning in the proposed model which helps to avoid over-fitting to training domains.
V Discussion
TABLE III: Comparison with other works for downstream datasets. # indicates the models where the downstream dataset is contained in the pre-training dataset. Results in bold indicate the best performing model, while those underlined indicate the second-best model. All numbers are weighted F1-scores computed over $5$ random initializations
Datasets WavLM [16] HuBERT [15] data2vec [32] emotion2vec [42] SONAR [22] SALMONN [20] CARE
Params:$94$M Params:$94$M Params:$94$M Params:$94$M Params:$600$M Params:$8$B Params:$14$B Params:$160$M
IEMOCAP-4 $65.85$ $64.96$ $62.68$ $67.45^{\#}$ $59.43$ $\mathbf{75.78}^{\#}$ $\underline{72.92}^{\#}$ $69.39$
IEMOCAP-6 $51.67$ $50.72$ $46.01$ $54.05^{\#}$ $43.48$ $\mathbf{59.31}^{\#}$ $\underline{58.07}^{\#}$ $55.02$
MELD $45.61$ $45.30$ $41.89$ $47.60^{\#}$ $43.16$ $\mathbf{53.34}$ $\underline{52.59}$ $48.05$
CMU-MOSI $64.09$ $62.53$ $59.67$ $66.45$ $\underline{74.61}$ $\mathbf{78.03}$ $72.80$ $66.74$
DAIC-WOZ $63.15$ $65.87$ $\underline{67.81}$ $61.62$ $64.32$ $62.61$ $64.72$ $\mathbf{68.49}$
RAVDESS-Song $50.45$ $\underline{53.46}$ $38.53$ $48.49$ $11.81$ $50.20$ $51.92$ $\mathbf{60.11}$
CaFE $66.58$ $66.53$ $48.75$ $59.27$ $5.73$ $59.86$ $\underline{69.84}$ $\mathbf{76.98}$
EmoDB $66.54$ $66.94$ $48.87$ $64.37$ $10.2$ $\underline{82.80}$ $82.15$ $\mathbf{83.41}$
Avg. $59.24$ $59.54$ $51.53$ $58.66$ $39.09$ $65.24$ $\underline{65.63}$ $\mathbf{66.02}$
V-A Comparison with prior works
A comparative analysis with several recent studies is presented in Table III. These baseline models are categorized into two groups based on the number of parameters used during inference: base models (parameter size $<200$M), and large models ($>500$M), which also include LLM based models. The following observations are made for each category:
V-A1 Base models
We compare HuBERT [15], WavLM [21], data2vec [32] and emotion2vec [42] representations as the baseline models in this category. Among these baseline systems, the emotion2vec is also pre-trained on IEMOCAP and MELD datasets, partially explaining the improved results seen on the downstream tasks on these datasets. While CARE performs similar to emotion2vec on CMU-MOSI, it achieves a significant improvement over all the base-sized models on other datasets. On the average, the proposed CARE achieves a relative improvement of $15.6$% over the best baseline model (HuBERT).
V-A2 Large models
SONAR [22] is selected as the speech encoder in this category. For the six English-based datasets, the pre-trained English speech encoder [^4] https://dl.fbaipublicfiles.com/SONAR/spenc.eng.pt is used, while the French [^5] https://dl.fbaipublicfiles.com/SONAR/spenc.v3ap.fra.pt and German [^6] https://dl.fbaipublicfiles.com/SONAR/spenc.v3ap.deu.pt speech encoders are utilized for the CaFE and EmoDB datasets, respectively. Similar to the CARE backend, the layer representations from the SONAR encoder are linearly combined and the classification head is trained on the downstream task. Although SONAR has nearly four times the parameter size of CARE, our proposed model outperforms SONAR across all datasets except CMU-MOSI.
V-A3 LLM based models
Two versions of SALMONN [20] ($8$B and $14$B) [^7] https://huggingface.co/tsinghua-ee/SALMONN are considered as examples of LLM-based models. A direct zero-shot evaluation of the LLM based models achieved poor results. Thus, the same framework used in CARE and other baseline models was followed for the LLM based evaluations as well. The internal representations from all layers ($41$ layers for SALMONN $14$B and $33$ layers for SALMONN $8$B) are aggregated using a convex combination, and the classification head (two-layer feed forward network) is trained for each downstream dataset. Similar to emotion2vec, SALMONN includes IEMOCAP in its pre-training, leading to superior performance on IEMOCAP-4 and IEMOCAP-6 compared to CARE. The larger model size and extensive pre-training data allows SALMONN to outperform CARE by $10\%$ and $34\%$ (relative improvements) on the MELD and CMU-MOSI datasets, respectively. However, on the remaining three tasks, CARE surpasses the SALMONN models, achieving relative improvements of $17\%$ and $24\%$ on the RAVDESS-song and CaFE datasets, respectively.
V-A4 Key takeaways
On average, CARE emerges as the top-performing model across the eight datasets. The SALMONN $13$B model, which has nearly $80$ times more parameters than CARE, is the next best-performing model. Although LLM-based models excel in emotion recognition for some of the in-domain tasks, their performance on the out-of-domain tasks suggests that they may not generalize well on diverse tasks and multi-lingual emotional speech.
TABLE IV: Results on the downstream datasets (weighted F1-score) with modifications in the semantic arm.
Dataset Method WF1
IEMOCAP(4-class) CARE-Trans $65.89$
CARE-FT $68.16$
CARE $\mathbf{69.39}$
IEMOCAP(6-class) CARE-Trans $51.19$
CARE-FT $52.37$
CARE $\mathbf{55.02}$
MELD CARE-Trans $46.57$
CARE-FT $47.93$
CARE $\mathbf{48.05}$
CMU-MOSI CARE-Trans $65.97$
CARE-FT $64.27$
CARE $\mathbf{66.74}$
DAIC-WOZ CARE-Trans $65.93$
CARE-FT $67.43$
CARE $\mathbf{68.49}$
TABLE V: Results on the downstream datasets (weighted F1-score) with different initializations of the acoustic and common arms.
Datasets Random init. HuBERT init. Data2vec init. WavLM init.
IEMOCAP-4 $66.72$ $67.65$ $66.76$ $\mathbf{69.39}$
IEMOCAP-6 $51.47$ $51.40$ $52.98$ $\mathbf{55.02}$
MELD $46.41$ $46.97$ $47.19$ $\mathbf{48.05}$
CMU-MOSI $65.07$ $66.26$ $\mathbf{68.14}$ $66.74$
DAIC-WOZ $67.56$ $65.19$ $66.59$ $\mathbf{68.49}$
RAVDESS-Song $57.81$ $\mathbf{60.30}$ $55.09$ $60.11$
CaFE $73.83$ $72.23$ $62.46$ $\mathbf{76.98}$
EmoDB $78.61$ $\mathbf{86.51}$ $77.19$ $83.41$
Avg. $63.44$ $64.56$ $62.05$ $66.02$
V-B Adapters in the semantic arm
The results in the previous sections used the semantic arm with pre-trained RoBERTa model and convolutional adapters. We make two modifications to this architecture:
  1. 1.
    CARE-Trans.: Removing the convolutional adapters while the RoBERTa transformer layers are updated in pre-training.
  2. 2.
    CARE-FT: Keeping the convolutional adapters, we update all the parameters (conv. adapters and transformer weights) in the semantic arm.
The results for these two modifications are shown in Table IV. The setup, where the transformers are also trained during pre-training, yields lower performance on the downstream datasets. The results also show that the proposed convolution layer based adaptation strategy offers a good compromise of speech representation learning capability from text-based models and parameter efficiency.
V-C Initialization of Acoustic Arm
As indicated in Section III, the common and the acoustic arm of CARE are initialized with the WavLM-base model weights. We present the results of our method when this initialization is modified to i) random initialization, ii) HuBERT-base [15] or iii) data2vec-base [32] (Table V). The HuBERT-base initialization is seen to outperform the WavLM-base version of CARE for the RAVDESS-song and EmoDB datasets. However, on the average, the WavLM initialization achieves the best performance.
V-D Choice of acoustic targets
We run an ablation experiment where the acoustic arm is trained with targets based on eGeMAPS [9] features extracted from the openSMILE toolkit [25]. The PASE+ targets of the acoustic arm of the CARE model is replaced by the eGeMAPS features. The performance of this model, called CARE (eGeMAPS), is shown in Fig. 2. The performance of the baseline model trained with eGeMAPS as input features is lower than that of the baseline with PASE+ features. Subsequently, the average performance of CARE (eGeMAPS) is lower than of CARE with PASE+ targets.
Figure 2: Performance of CARE when different acoustic targets are used. The model with eGeMAPS as features is trained similarly to that of the PASE+ baseline. All numbers are shown as the average of $5$ random initializations.
V-E Continued pre-training of WavLM
To explore the impact of pre-training setup in the proposed CARE, we continued the pre-training of the publicly available WavLM-base model, using the MSP-PODCAST dataset. This was done following the WavLM pre-training procedure, with masked language modeling loss, for an additional $200,000$ steps (similar to the CARE). The results of this experiment are shown in Fig. 3, wherein the performance of the continually pre-trained WavLM model is denoted by WavLM-Cont. Pretrain. The continued pre-training improves the performance of the WavLM-base on some of the downstream tasks like IEMOCAP. However, with the exception of IEMOCAP dataset, the WavLM system’s performance is worse than that of the acoustic arm of the CARE model. As all these models have the same size ($94$M), this ablation highlights the benefits of distillation with the PASE+ representations.
Figure 3: Comparison of the performance when WavLM is continually pre-trained on MSP-PODCAST. The performance of the acoustic arm of CARE is shown for reference. Here, RAVDESS refers to the RAVDESS-Song dataset.
TABLE VI: Results for the MSP-IMPROV dataset. CCC stands for Concordance Correlation Coefficient while V, A, D stand for valence, arousal and dominance respectively.
Method CCC-V CCC-A CCC-D
WavLM-base [16] $0.51$ $0.64$ $0.47$
emotion2vec [42] $0.5$ $0.61$ $0.49$
SALMONN-$8$B [20] $0.53$ $\mathbf{0.67}$ $\underline{0.52}$
SALMONN-$14$B [20] $\underline{0.56}$ $0.65$ $0.51$
CARE $\mathbf{0.57}$ $\underline{0.66}$ $\mathbf{0.53}$
V-F Emotional attribute prediction
The emotion recognition can be posed as a regression problem, where valence, arousal and dominance of a particular utterance is predicted [69]. We use the MSP-IMPROV [70] for this purpose. This is an audio-visual dataset that consists of $12$ actors eliciting a set of sentences in different emotions. The dataset consists of $8438$ utterances with valence, arousal and dominance values (ranging from $1$ to $5$). We first split the dataset in $12$ parts, where each part contains utterances corresponding to $10$ training speakers, while speech from the two other speakers are used for validating and testing the model. The performance is measured as the average over these $12$ parts.
We use the concordance correlation coefficient (CCC) as the metric, which indicates the agreement between the predicted and the ground truth attribute scores for this dataset. Denoting the mean, variance of ground truth by $\mu_{g}$, $\sigma_{g}^{2}$ and predicted scores by $\mu_{p}$, $\sigma_{p}^{2}$, the CCC is defined as
$CCC=\frac{2\rho\sigma_{g}\sigma_{p}}{\sigma_{p}^{2}+\sigma_{g}^{2}+(\mu_{g}-\mu_{p})^{2}}$ (4)
In Eq. 4, $\rho$ is the Pearson’s correlation coefficient between the ground truth and the predicted scores. For training the downstream model, the objective is to increase the CCC between the ground truth and the predicted values for each of the dimensions of valence, arousal and dominance. The results for this dataset along with other baseline models are shown in Table.VI. We note that for this task, the CARE embeddings achieve the best results in terms of the valence and dominance attributes, while the performance on arousal is marginally better for the SALMONN-$8$B model.
TABLE VII: Results on the different datasets for L1111 and L2222 loss
Dataset Loss WF1
IEMOCAP(4-class) L1 $68.54$
L2 $\mathbf{69.39}$
IEMOCAP(6-class) L1 $54.82$
L2 $\mathbf{55.02}$
MELD L1 $47.23$
L2 $\mathbf{48.05}$
CMU-MOSI L1 $66.61$
L2 $\mathbf{66.74}$
DAIC-WOZ L1 $67.53$
L2 $\mathbf{68.49}$
V-G Choice of loss function
We present the results on the different downstream datasets by modifying the loss function to mean absolute loss (L1111) instead of L2222 loss (Table VII). In most cases, the model trained with L2222 loss fares better than the one trained with L1111 loss.
Figure 4: Distribution of cosine similarities for different layer representations for the acoustic and the semantic arm. The top three plots are when two speech signals belonging to different speakers, same emotion and same content are processed by CARE. The bottom three plots refer to the setting with different speakers, same emotion and different content.
V-H Visualization of layers of CARE
In order to interpret the pre-trained model representations learnt by the acoustic and semantic arm of CARE, we probe the representations from the transformer blocks of each arm. We use the English part of the Emotional Speech Dataset (ESD) [71] for this analysis. We form pairs of utterances, where both the utterances of a pair have the same emotional label in all cases and they are derived from two different speakers. A total of $1750$ pairs are considered and the cosine similarities of the pooled representations (for $3$ transformer layers) are shown in Fig.4. The two panels indicate the settings where the speech content in the two utterances are the same (top-panel) or different (bottom-panel). Note that, the representations are obtained from the pre-trained model with unsupervised data, hence we do not expect the model to fully encode emotions. Following are the observations from this visualization:
  1. 1.
    The semantic arm in layer $11$ has slightly more similarity than the acoustic arm for the set of pairs with same content (top-panel).
  2. 2.
    When the spoken content is different, the acoustic arm has significantly higher similarity than the semantic arm, indicating that the acoustic arm is beneficial when the emotion information cannot be reliably predicted from the textual content of the audio.
VI Summary
Key Highlights: In this paper, a pre-training technique for content and acoustic encoding of emotional speech is provided. The proposed architecture, termed CARE, learns an enriched representation of acoustic and semantic information. The acoustic arm uses supervision from low-level descriptors of speech, while the semantic arm is distilled using text representations of the speech transcripts. We also propose a novel adaptation strategy for text-based models in speech representation learning using convolutional neural network layers. The CARE model, with experiments on $8$ downstream tasks, is seen to significantly outperform models of comparable sizes on most of the datasets. Further, the CARE is also observed to generalize better than LLM based models with large parameter sizes. The importance of the different components of the proposed model, along with the different design choices, are established through ablation studies.
Limitations and future scope: The MSP-PODCAST dataset is used for pre-training CARE, which has only $230$ hours of emotional speech data. Training the model on a bigger corpus may improve the performance of the model further. Another limitation of this work, is the relatively lower performance on some in-domain speech datasets compared to LLM-based models, like the CMU-MOSI dataset. In future, we plan to extend the CARE approach to multi-modal speech-text emotion recognition tasks. Further, other masked losses will be explored instead of the L2 based distillation loss utilized in the current work.